Robert is 3 times as older as his son; but 14 years ago he was 10 times as older as his son. Find the age of his son.
Answer by stanbon(75887)

Robert is 3 times as old as his son; but 14 years ago he was 10 times as older as his son. Find the age of his son.
---------------------
Equations:
r = 3s
r-14 = 10(s-14)
---
Substitute for "r" and solve for "s":
3s-14 = 10s-140
7s = 126
s = 18 (son's age now)
========================
Cheers,
Stan H.